# Motion Before Action: Diffusing Object Motion as Manipulation Condition

### v1

@@ V-D Open Drawer
In this experiment, we account for the fact that delicate operations can be easily influenced by scene setup, particularly issues like point cloud occlusion when the robotic arm manipulates drawers at low heights. To minimize the impact of randomness, we doubled the number of test trials from the original plan, increasing it to 40 trials. Our test positions cover a variety of positions and orientations in the workspace, considering both fully closed and partially open states of the drawer. We also include scenarios where the initial drawer position causes point cloud occlusion, ensuring a comprehensive evaluation of the policies’ robustness.
We report the average success rates of the two task stages (Grasp: grasp the handle; Pull: pull the drawer until open) as evaluation metrics to accurately assess the execution capabilities of the policies.
Method Success Rates (%) $\uparrow$
Grasp Pull
RISE $37.5$ $37.5$
RISE w. MBA $52.5$ $52.5$
TABLE V: Experimental Results of the Open Drawer task.
@@
The experimental results shown in Tab. V indicate that RISE with MBA outperforms the baseline RISE in both stages of the task. This aligns with our findings in simulation experiments, further confirming that integrating MBA significantly enhances the baseline policy in tasks that require precise object localization and fine manipulation.
VI Conclusion

### v3

@@ V-D Open Drawer
In this experiment, we account for the fact that delicate operations can be easily influenced by scene setup, particularly issues like point cloud occlusion when the robotic arm manipulates drawers at low heights. To minimize the impact of randomness, we double the number of test trials from the original plan, increasing it to 40 trials. Our test positions cover a variety of positions and orientations in the workspace, including areas within both the inner circle and the outer ring of the workspace plane. Each area receives 20 test trials. We also include scenarios where the initial drawer position causes point cloud occlusion, ensuring a comprehensive evaluation of the policies’ robustness. We report the average success rates over all the trials.
@@
The experimental results are shown in Table. III indicate that policies with MBA outperform the baselines in both stages of the task. This aligns with our findings in simulation experiments, further confirming that integrating MBA significantly enhances the policies in tasks that require precise object localization and fine manipulation.
V-E Pour Balls
The Pour Balls task is a challenging one, where the objective is to lift a cup filled with 10 balls and pour them into a bowl. The difficulty arises from two main factors:
  1. 1)
    The cup is a non-cubic object, with varying width at different heights. Therefore, the gripper must learn a precise visual-to-motion control policy to grasp the cup at the correct height. If the gripper is too wide, it will fail to hold the cup, causing it to fall; if the gripper is too narrow, it will knock over the cup.
  2. 2)
    This is a 6-DoF task, and during the pouring process, if the translation and rotation are not properly controlled, or the gripper’s force is not correctly adjusted, the cup will either rotate or fall, resulting in task failure.
In this round of experiments, we still randomize the object positions. We conduct 20 trials. In this experiment, we consider three evaluation metrics: the average success rate of pouring balls into the bowl, the average number of balls poured into the bowl across all trials, and the average success rate of picking up the cup.
In this experiment, MBA achieves more than 15% the success rate compared to RISE and 30% compared to DP, as shown in Table III. Notably, under the same successful conditions of picking up the cup, MBA is also able to pour the balls with greater precision. This indicates that MBA can accurately capture the relationship between object pose variations and corresponding actions in 6-DoF tasks, making it well-suited for handling intricate manipulation scenarios.
V-F Comparative Experiment to Flow-based methods
In this section, we compare MBA against the representative flow-based method ATM [ATM] to demonstrate that the pose prediction paradigm yields higher quality action generation for the policy compared to prediction in the visual space. Specifically, both ATM and MBA employ a diffusion action head for prediction; MBA conditions on predicted future object pose sequences, whereas ATM conditions on keypoint flow predicted by a Track Transformer.
We evaluate MBA on three 2D policy-fit tasks and found that it outperforms ATM across most metrics, as shown in Table IV. Notably, in tasks requiring fine-grained manipulation, such as Open Drawer, the flow-based approach’s tracking is confined to a limited number of pixels in the visual space, making it challenging to capture the pose of the handle, which can lead to grasping failures. In contrast, tasks with higher degrees of freedom, like Pour Balls, necessitate more comprehensive spatial motion modeling and precise control over the object’s rotational dynamics. In these scenarios, flow-based policies appear to be at a disadvantage. The results validate our hypothesis regarding the “vision-motion” gap, underscoring the necessity of modeling pose information.
Task Metric DP w. ATM [ATM] DP w. MBA
Put Bread into Pot Success (%) $\uparrow$ $40$ $45$
Attempt $\downarrow$ $2$ $16$
Open Drawer Success (%) $\uparrow$ $5$ $30$
Pour Balls Pour (%) $25$ $40$
Balls $\uparrow$ $2.05$ $3.60$
Pick (%) $35$ $60$
TABLE IV: Comparison of policy models under different object motion conditions: MBA (ours): $6$D poses vs. ATM: 2D point flow.
V-G Inference Speed
We evaluated the average inference time (excluding visual backbone processing time for experimental consistency across different backbones) for DP, ATM, and DP with MBA. DP achieved 95.98 ms, ATM 105.85 ms, while DP with MBA required 197.50 ms. This demonstrates the increased computational cost associated with MBA’s precise control. ATM benefits from its transformer-based flow tracking, avoiding computationally expensive denoising steps.
VI Limitations
MBA still exhibits several limitations. First, its inference efficiency can be improved. Future work might focus on replacing the existing denoising paradigm with more computationally efficient models [ACT, su2025dense]. Second, MBA’s object pose supervision relies on ground-truth data from a MoCap system, which incurs high acquisition costs. Borrowing from flow-based methods that utilize flow generative models to provide ground truth [ATM, YuanGeneralFlow] offers a promising alternative. Correspondingly, leveraging foundation 6D pose estimation models [wen2024foundationpose, hirschorn2023pose] for data annotation is also a valuable direction. Third, MBA is limited in handling varying numbers of objects, as its predefined observation vector constrains it to a fixed object count. Finally, the manipulation of deformable objects, whose 6D pose is not trackable, remains unaddressed. These limitations point to important directions for future research.
VII Conclusion